OpenStudy (curry):

s=5m r= 10m find theta

OpenStudy (anonymous):

theta??? where is theta?? I only see s, r and m

OpenStudy (curry):

s/r = tehta the answer is 60 degress for theta

OpenStudy (curry):

but if u do s/r it doesnt work

OpenStudy (anonymous):

ohhhh i see lol

OpenStudy (anonymous):

Theta in radians is product of arclength (s) and radius (r)

OpenStudy (anonymous):

REMEMBER: RADIANS!

OpenStudy (curry):

s=(pi(r(theta)/360

OpenStudy (curry):

so how do u solve this i keep getting 28

OpenStudy (curry):

but it is supposed to be 60

OpenStudy (anonymous):

wouldn't it be 2pi/360?

OpenStudy (curry):

nope

OpenStudy (anonymous):

60 degrees is incorrect

OpenStudy (anonymous):

http://www.mathopenref.com/arclength.html that shows it as being 29

OpenStudy (anonymous):

which is what i got through calculation.

OpenStudy (anonymous):

5*360/(2pi*10)=theta